OpenStudy (cutestgirlaround):

5/9s+13/9s=24 I think s is -6 am I right

OpenStudy (cutestgirlaround):

@Awolflover1

Awolflover1 (awolflover1):

-6 is incorrect

OpenStudy (cutestgirlaround):

oh well I know if 5/9 + 13/9 = 18/9 so I thought I would do 18-24 which = -6

Awolflover1 (awolflover1):

you are correct but 2s=24 then Divide both sides by 2

OpenStudy (cutestgirlaround):

oh yeah because 5/9 + 13/9 = 18/9 which = 2 right?

Awolflover1 (awolflover1):

yes \[24\div2\]